# Alignment-Sensitive Minimax Rates for Spectral Algorithms with Learned Kernels

### v1

@@ 3 Effective Span Dimension and Span Profile
Principal Component (PC): $\psi_{\nu}^{\operatorname{PC}}(\lambda)=1_{\{\lambda<\nu\}},$ $\widehat{\theta}_{j}^{\operatorname{PC},\nu}$ $=1_{\{\lambda_{j}\geq\nu\}}\,z_{j}.$ (7)
@@ 7 Extension to Correlated Noise and Fixed-design Linear Model
This section extends the concepts of ESD and span profile, developed in Section˜3 for the sequence model, to the setting of fixed design linear regression. In addition, we demonstrate how the minimax optimal prediction risk in this setting can be characterized using the span profile, paralleling the analysis in Section˜4.
@@ 7.1 Strategy of reduction
Before introducing the linear model, it is helpful to outline our general transformation strategy in the context of a sequence model with correlated noise. Suppose $d\in\mathbb{N}_{+}$ and the observations is
@@
where $\mathbf{\Sigma}_{\xi}\in\mathbb{R}^{d\times d}$ is known, symmetric, and positive definite. For correlated sequence model, it is usually of interest to measure the estimation error using the squared Mahalanobis distance defined as $L(\bm{\theta}^{\star},\widehat{\bm{\theta}}^{\star})=(\widehat{\bm{\theta}}^{\star}-\bm{\theta}^{\star})^{\top}\mathbf{\Sigma}_{\xi}^{-1}(\widehat{\bm{\theta}}^{\star}-\bm{\theta}^{\star})$.
@@
$d_{\mathbf{\Sigma}_{\xi}}^{\dagger}\left(\tau;\bm{\theta}^{\star}\right):=d^{\dagger}\left(\tau;\bm{\tilde{\theta}}^{\star}=\mathbf{L}\bm{\theta}^{\star}\right)$
@@
so the minimax risk in terms of $\mathbb{E}\|\bm{\theta}^{\star}-\widehat{\bm{\theta}}^{\star}\|^{2}$ can still be characterized sharply when the condition number of $\mathbf{\Sigma}_{\xi})$ is bounded.
This strategy of reducing a complex model to a simple model will be used in our analysis of linear model and also the RKHS regression in Section˜8.
@@ 7.2 Linear Model
Consider the following fixed design linear regression model:
@@
where $\mathbf{Y}\in\mathbb{R}^{n}$ is the vector of observations, $\mathbf{X}\in\mathbb{R}^{n\times p}$ is the fixed design matrix of rank $r\leq\min(n,p)$, $\bm{\beta}^{*}\in\mathbb{R}^{p}$ is the unknown vector of true coefficients, and $\bm{\epsilon}\in\mathbb{R}^{n}$ is the noise vector. We assume the components of $\bm{\epsilon}$ are uncorrelated with mean zero and variance $\sigma_{0}^{2}$. For this model, we consider the loss (in-sample prediction error) $\mathcal{L}(\widehat{\bm{\beta}};\bm{\beta}^{*})=\frac{1}{n}\|\mathbf{X}\left(\widehat{\bm{\beta}}-\bm{\beta}^{*}\right)\|^{2}$ and risk $\mathcal{R}(\widehat{\bm{\beta}};\bm{\beta}^{*})=\mathbb{E}\mathcal{L}(\widehat{\bm{\beta}};\bm{\beta}^{*})$. For random design linear regression, we treat it as a special case of RKHS regression and discuss it in Section˜8.
@@
Multiplying the model Equation˜15 by $\frac{1}{\sqrt{n}}\mathbf{U}_{\cdot,[r]}^{\top}$, we obtain a $r$-dimensional transformed model:
@@
For any estimator $\widehat{\bm{\beta}}$ for the linear model Equation˜15, define $\widehat{\bm{\theta}}=\mathbf{S}_{[r],\cdot}\mathbf{V}^{\top}\widehat{\bm{\beta}}$. We can then write the prediction risk as $n^{-1}\mathbb{E}\|\mathbf{X}\widehat{\bm{\beta}}-\mathbf{X}\bm{\beta}^{*}\|^{2}=\mathbb{E}\|\mathbf{U}\mathbf{S}\mathbf{V}^{\top}\widehat{\bm{\beta}}-\mathbf{U}\mathbf{S}\mathbf{V}^{\top}\bm{\beta}^{*}\|^{2}=\mathbb{E}\|\widehat{\bm{\theta}}-\bm{\theta}^{*}\|^{2}$. Conversely, given an estimator $\widehat{\bm{\theta}}$ for the sequence model Equation˜17, we can define $\widetilde{\bm{\beta}}=\mathbf{V}\mathbf{S}^{\dagger}\widehat{\bm{\theta}}$ where $\mathbf{S}^{\dagger}\in\mathbb{R}^{p\times r}$ is an diagonal matrix whose diagonal elements are $\{1/s_{j}\}_{j\in[r]}$. It is easy to check that $\mathbf{S}_{[r],\cdot}\mathbf{V}^{\top}\widetilde{\bm{\beta}}=\widehat{\bm{\theta}}$. Therefore, we establish an equivalence between the model Equation˜17 and the model Equation˜15.
@@ 7.3 ESD for linear models
The Principal Component Regression (PCR) estimator for $\bm{\beta}^{*}$ corresponds to the Principal Component (PC) estimator in the transformed space Equation˜17. Specifically, for any $k\in[r]$, define $\widehat{\bm{\beta}}^{\operatorname{PC},k}=\frac{1}{\sqrt{n}}\mathbf{V}\mathbf{S}_{k}^{\dagger}\mathbf{U}_{\cdot,[r]}^{\top}\mathbf{Y}$, where $\mathbf{S}_{k}^{\dagger}\in\mathbb{R}^{p\times r}$ is a diagonal matrix whose diagonal elements are $\{\frac{1}{s_{j}}1_{\{s_{j}\geq s_{\pi_{k}}\}}\}$. In the $\mathbf{Z}$ space, this means
@@
$\mathcal{B}_{K}^{(n)}=\left\{\bm{\beta}\in\mathbb{R}^{p}:d^{\dagger}(\sigma_{0}^{2}/n;\bm{\beta}^{*},\mathbf{X})\leq K\right\},$ (19)
@@
Through this extension, the span profile framework connects the optimal prediction performance in fixed design linear regression to the alignment between the signal structure (transformed via the design matrix) and the spectrum derived from the design matrix’s singular values.
@@ 7.4 Numerical illustration
We are interested in the ESD and the minimum risk for different transformation of the design matrix. For this purpose, we introduce a class of non-orthogonal column transformations indexed by $\alpha>0$ as follows:
@@
The transformed design $\mathbf{X}(\alpha)=\mathbf{X}_{0}\mathbf{A}(\alpha)$ and the correspondingly transformed coefficient vector is $\bm{\beta}(\alpha)=\mathbf{A}(\alpha)^{-1}\bm{\beta}^{*}$. This family of transformation will change the order of spectrum, from well-aligned to misaligned. We are interested in the following at each $\alpha$:
@@
Figure 3 plots $d^{\dagger}(\alpha)$ (solid) and the rescaled oracle risk defined as $n\mathcal{R}_{*}(\alpha)/\sigma_{0}^{2}$ (dashed) against $\alpha$. The two curves coincide over the entire path, which empirically verifies the bound in Proposition˜7.2 that $\mathcal{R}_{*}(\alpha)\asymp\frac{\sigma_{0}^{2}}{n}\,d^{\dagger}(\alpha)$. As $\alpha$ grows, the diagonal stretch $\mathbf{A}(\alpha)$ shifts signal energy towards directions that carry smaller singular values. This raises $d^{\dagger}$ as well as the achievable risk increases.
This experiment illustrate that ESD, rather than the raw spectrum decay, is the pivotal measure that fully governs learnability.
@@ 8 Extension to RKHS Regression
This section extends the concepts of Effective Span Dimension (ESD) and span profile, developed in Section˜3, to the setting of RKHS regression. Here we will focus on the simple case where the eigenfunctions of the kernel are fully known and computable and leave a thorough analysis in future studies.
@@ 8.1 RKHS regression
$y_{i}=f^{*}(x_{i})+\epsilon_{i},\quad\epsilon_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}(0,\sigma_{0}^{2}),\quad i=1,\dots,n,$ (20)
where $x_{i}\sim\mu$ i.i.d., and $f^{*}\in L^{2}(\mathcal{X},\mu)$ is the target function. We use a kernel $\mathbf{k}(\cdot,\cdot)$ with Mercer decomposition $\mathbf{k}(x,x^{\prime})=\sum_{j=1}^{\infty}\lambda_{j}\psi_{j}(x)\psi_{j}(x^{\prime})$, where $\{\psi_{j}\}_{j=1}^{\infty}$ form an orthonormal basis for $L^{2}(\mathcal{X},\mu)$ and $\bm{\lambda}=\{\lambda_{j}\}_{j\geq 1}$ is the sequence of eigenvalues not necessarily sorted. For simplicity, we assume there is no ties among the eigenvalues and let $\pi$ be the permutation sorting eigenvalues $\lambda_{j}$ in descending order so that $\lambda_{\pi_{1}}>\lambda_{\pi_{2}}>\dots>\lambda_{\pi_{k}}>\ldots$. The coefficients of $f^{*}$ in this basis are $\theta_{j}^{*}=\langle f^{*},\psi_{j}\rangle_{L^{2}(\mu)}$. An estimator $\hat{f}$ has risk $\mathcal{R}(\hat{f};f^{*})=\mathbb{E}\|\hat{f}-f^{*}\|_{L^{2}(\mu)}^{2}$. If we define $\hat{\theta}_{j}=\langle\hat{f},\psi_{j}\rangle_{L^{2}(\mu)}$, we can also write the risk as
@@
Here we kindly remind the reader that in RKHS, we use to the subscript $i$ to index samples while using $j$ to index eigen-coordinates. The subscript $j$ aligns with the use of notation in the sequence model, where we use indices $j$ to denote coordinates.
@@
Before we introduce the definition of ESD, we demonstrate that the approximation in Equation˜3 can be made exact by increasing the variance to encompass the approximation error.
@@
$=\sum_{k\geq 1}\theta_{k}^{*}\underbrace{\left(\frac{1}{n}\sum_{i=1}^{n}\psi_{k}(x_{i})\psi_{j}(x_{i})\right)}_{:=G_{kj}}\;+\;\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}\,\psi_{j}(x_{i})$
@@
Since $x_{i}\stackrel{{\scriptstyle iid}}{{\sim}}\mu$ and $\{\psi_{j}\}$ is orthonormal in $L^{2}(\mathcal{X},\mu)$, then $\mathbb{E}[G_{kj}]=\mathbf{1}_{\{k=j\}}$. Hence $\mathbb{E}[z_{j}]=\theta_{j}^{*}$.
@@
where we have defined $\Delta_{j}=(G_{jj}-1)\theta_{j}^{*}+\left(\sum_{k\geq 1,k\neq j}G_{kj}\theta_{k}^{*}\right)$. This term does not appear in the sequence model, and its randomness purely comes from random covariate $x_{i}$’s. as $n\to\infty$, this term vanishes because $G_{jj}=n^{-1}\sum_{i}\psi_{j}\left(x_{i}\right)^{2}\to 1$ and $G_{kj}=n^{-1}\sum_{i=1}^{n}\psi_{j}\left(x_{i}\right)\psi_{j^{\prime}}\left(x_{i}\right)\to 0$. Furthermore, since $\mathbb{E}(\xi_{j}\mid\{x_{i}\}_{i\in[n]})=0$, we have
@@
The presence of $\Delta_{j}$ effectively inflates the variance in $z_{j}$ to $\sigma_{0}^{2}/n+\mathrm{Var}(\Delta_{j})$. One can show that $\mathrm{Var}(\Delta_{j})=\mathrm{Var}(f^{*}(x)\psi_{j}(x))$, which is bounded by $\|f\|_{\infty}^{2}$. This is how we will control the impact of $\Delta_{j}$ in the following development.
@@ 8.2 ESD for RKHS Regression
We start with analyzing the counterpart of the PC estimator, the Kernel Principal Component Projection Estimator (KPCPE), defined as
$\hat{f}^{\operatorname{PC},k}(x):=\sum_{j:\lambda_{j}\geq\lambda_{\pi_{k}}}z_{j}\psi_{j}(x),$ (22)
@@
The risk $\mathcal{R}_{k}:=\mathbb{E}\|\hat{f}_{k}^{\text{KPCPE}}-f^{*}\|_{L^{2}(\mu)}^{2}$ decomposes into squared bias $B(k)$ and variance $V(k)$. Since $\mathbb{E}[\psi_{j^{\prime}}(x_{i})\psi_{j}(x_{i})]=1_{\{j=j^{\prime}\}}$, we have $\mathbb{E}[z_{j}]=\theta_{j}^{*}$. Therefore, the bias is due to truncation as
@@
Under Assumption 8.1, $\tau_{j}^{2}\leq\mathbb{E}[f^{*}(x)^{2}\psi_{j}(x)^{2}]\leq\|f^{*}\|_{\infty}^{2}$. Subsequently, the variance is bounded by $V(k)\leq\frac{k}{n}(\sigma_{0}^{2}+\|f^{*}\|_{\infty}^{2})$. This motivates use to define the effective noise variance per component as
@@
The kernel $\mathbf{k}$ is said to be $(K,n)$-regular if there are some constants $c_{1}<1$ and $C_{1}$ such that $\sum_{i\leq c_{1}K}\lambda_{\pi_{i}}^{-1}\leq C_{1}n$.
@@
$\inf_{\hat{f}}\sup_{f^{*}\in\mathcal{F}_{K,\mathbf{k}}^{(n)}}\mathcal{R}(\hat{f};f^{*})\asymp\frac{\bar{\sigma}^{2}K}{n},$ (29)
@@
Combining Theorem˜8.5 and Proposition˜8.3, the optimally tuned KPCPE estimator is minimax rate optimal over $\mathcal{F}_{K,\mathbf{k}}^{(n)}$, with rate $\bar{\sigma}_{0}^{2}K/n$.
The KPCPE serves as a simple benchmark for spectral methods. This analysis, via the ESD, characterizes the performance of the optimally tuned KPCPE based directly on the properties of the specific signal $f^{*}$ (via $\theta^{*}$) and kernel spectrum $\lambda$, without requiring standard assumptions like source conditions or polynomial eigenvalue decay. Therefore, we consider the ESD evaluated at the design-adjusted noise level $\sigma^{2}$ as a key measure of statistical complexity in RKHS regression. In summary, the span profile framework provides a unified perspective on generalization performance of spectral methods on a variety of models.
@@
Following the framework in Section˜4, we can quantify a class of populations using a quota sequence $\mathbf{K}=\{K_{n}\}_{n=1}^{\infty}$. For some $n_{0}\in\mathbb{N}_{+}$, define
$\mathcal{F}_{\mathbf{K},\mathbf{k}}=\{f^{*}\in L^{2}(\mathcal{X},\mu)\cap L^{\infty}(\mathcal{X},\mu):\|f^{*}\|_{\infty}\leq\sigma_{0}C_{0},\quad d^{\dagger}(\bar{\sigma}^{2}/n;f^{*},\mathbf{k})\leq K_{n},\forall n\geq n_{0}\},$ (30)
@@
$\inf_{\hat{f}}\sup_{f^{*}\in\mathcal{F}_{\mathbf{K},\mathbf{k}}}\mathcal{R}(\hat{f},f^{*}).$ (31)
@@
$\inf_{\hat{f}}\sup_{f^{*}\in\mathcal{F}_{\mathbf{K},\mathbf{k}}}\mathcal{R}(\hat{f},f^{*})\asymp\bar{\sigma}^{2}\frac{K_{n}}{n}.$

### v2

@@ 3 Effective Span Dimension and Span Profile
Principal Component (PC): $\psi_{\nu}^{\operatorname{PC}}(\lambda)=\mathbf{1}_{\{\lambda<\nu\}},$ $\widehat{\theta}_{j}^{\operatorname{PC},\nu}$ $=\mathbf{1}_{\{\lambda_{j}\geq\nu\}}\,z_{j}.$ (7)
@@ 7 Extension to Correlated Noise and Fixed-Design Linear Model
This section extends the concepts of ESD and span profile, developed in Section˜3 for the sequence model, to the setting of fixed-design linear regression. In addition, we demonstrate how the minimax optimal prediction risk in this setting can be characterized using the span profile, paralleling the analysis in Section˜4.
@@ 7.1 Strategy of reduction
Before introducing the linear model, it is helpful to outline our general transformation strategy in the context of a sequence model with correlated noise. Suppose $d\in\mathbb{N}_{+}$ and the observations are
@@
where $\mathbf{\Sigma}_{\xi}\in\mathbb{R}^{d\times d}$ is known, symmetric, and positive definite. For a correlated sequence model, it is usually of interest to measure the estimation error using the squared Mahalanobis distance defined as $L(\bm{\theta}^{\star},\widehat{\bm{\theta}}^{\star})=(\widehat{\bm{\theta}}^{\star}-\bm{\theta}^{\star})^{\top}\mathbf{\Sigma}_{\xi}^{-1}(\widehat{\bm{\theta}}^{\star}-\bm{\theta}^{\star})$.
@@
$d_{\mathbf{\Sigma}_{\xi}}^{\dagger}\left(\tau;\bm{\theta}^{\star}\right):=d^{\dagger}\left(\tau;\bm{\tilde{\theta}}^{\star}=\mathbf{L}\bm{\theta}^{\star}\right).$
@@
so the minimax risk in terms of $\mathbb{E}\|\bm{\theta}^{\star}-\widehat{\bm{\theta}}^{\star}\|^{2}$ can still be characterized sharply when the condition number of $\mathbf{\Sigma}_{\xi}$ is bounded.
This strategy of reducing a complex model to a simple model will be used in our analysis of the linear model and in RKHS regression in Section˜8.
@@ 7.2 Linear model
Consider the following fixed-design linear regression model:
@@
where $\mathbf{Y}\in\mathbb{R}^{n}$ is the vector of observations, $\mathbf{X}\in\mathbb{R}^{n\times p}$ is the fixed-design matrix of rank $r\leq\min(n,p)$, $\bm{\beta}^{*}\in\mathbb{R}^{p}$ is the unknown vector of true coefficients, and $\bm{\epsilon}\in\mathbb{R}^{n}$ is the noise vector. We assume the components of $\bm{\epsilon}$ are uncorrelated with mean zero and variance $\sigma_{0}^{2}$. For this model, we consider the loss (in-sample prediction error) $\mathcal{L}(\widehat{\bm{\beta}};\bm{\beta}^{*})=\frac{1}{n}\|\mathbf{X}\left(\widehat{\bm{\beta}}-\bm{\beta}^{*}\right)\|^{2}$ and risk $\mathcal{R}(\widehat{\bm{\beta}};\bm{\beta}^{*})=\mathbb{E}\mathcal{L}(\widehat{\bm{\beta}};\bm{\beta}^{*})$. For random design linear regression, we treat it as a special case of RKHS regression and discuss it in Section˜8.
@@
Multiplying the model in Equation˜15 by $\frac{1}{\sqrt{n}}\mathbf{U}_{\cdot,[r]}^{\top}$, we obtain a $r$-dimensional transformed model:
@@
For any estimator $\widehat{\bm{\beta}}$ for the linear model Equation˜15, define $\widehat{\bm{\theta}}=\mathbf{S}_{[r],\cdot}\mathbf{V}^{\top}\widehat{\bm{\beta}}$. We can then write the prediction risk as $n^{-1}\mathbb{E}\|\mathbf{X}\widehat{\bm{\beta}}-\mathbf{X}\bm{\beta}^{*}\|^{2}=\mathbb{E}\|\mathbf{U}\mathbf{S}\mathbf{V}^{\top}\widehat{\bm{\beta}}-\mathbf{U}\mathbf{S}\mathbf{V}^{\top}\bm{\beta}^{*}\|^{2}=\mathbb{E}\|\widehat{\bm{\theta}}-\bm{\theta}^{*}\|^{2}$. Conversely, given an estimator $\widehat{\bm{\theta}}$ for the sequence model Equation˜17, we can define $\widetilde{\bm{\beta}}=\mathbf{V}\mathbf{S}^{\dagger}\widehat{\bm{\theta}}$ where $\mathbf{S}^{\dagger}\in\mathbb{R}^{p\times r}$ is a diagonal matrix whose diagonal elements are $\{1/s_{j}\}_{j\in[r]}$. It is easy to check that $\mathbf{S}_{[r],\cdot}\mathbf{V}^{\top}\widetilde{\bm{\beta}}=\widehat{\bm{\theta}}$. Therefore, we establish an equivalence between the model Equation˜17 and the model Equation˜15.
@@ 7.3 ESD for linear models
The Principal Component Regression (PCR) estimator for $\bm{\beta}^{*}$ corresponds to the Principal Component (PC) estimator in the transformed space Equation˜17. Specifically, for any $k\in[r]$, define $\widehat{\bm{\beta}}^{\operatorname{PC},k}=\frac{1}{\sqrt{n}}\mathbf{V}\mathbf{S}_{k}^{\dagger}\mathbf{U}_{\cdot,[r]}^{\top}\mathbf{Y}$, where $\mathbf{S}_{k}^{\dagger}\in\mathbb{R}^{p\times r}$ is a diagonal matrix whose diagonal elements are $\{\frac{1}{s_{j}}\mathbf{1}_{\{s_{j}\geq s_{\pi_{k}}\}}\}$. In the $\mathbf{Z}$ space, this means
@@
$\mathcal{B}_{K}^{(n)}=\left\{\bm{\beta}^{*}\in\mathbb{R}^{p}:d^{\dagger}(\sigma_{0}^{2}/n;\bm{\beta}^{*},\mathbf{X})\leq K\right\},$ (19)
@@
Through this extension, the span profile framework connects the optimal prediction performance in fixed-design linear regression to the alignment between the signal structure (transformed via the design matrix) and the spectrum derived from the design matrix’s singular values.
@@ 7.4 Numerical illustration
We are interested in the ESD and the minimum risk for different transformations of the design matrix. For this purpose, we introduce a class of non-orthogonal column transformations indexed by $\alpha>0$ as follows:
@@
The transformed design $\mathbf{X}(\alpha)=\mathbf{X}_{0}\mathbf{A}(\alpha)$ and the correspondingly transformed coefficient vector is $\bm{\beta}(\alpha)=\mathbf{A}(\alpha)^{-1}\bm{\beta}^{*}$. These transformations will change the ordering of the spectrum, from well-aligned to misaligned. We are interested in the following at each $\alpha$:
@@
Figure 3 plots $d^{\dagger}(\alpha)$ (solid) and the rescaled oracle risk defined as $n\mathcal{R}_{*}(\alpha)/\sigma_{0}^{2}$ (dashed) against $\alpha$. The two curves coincide over the entire path, which empirically verifies the bound in Proposition˜7.2 that $\mathcal{R}_{*}(\alpha)\asymp\frac{\sigma_{0}^{2}}{n}\,d^{\dagger}(\alpha)$. As $\alpha$ grows, the diagonal stretch $\mathbf{A}(\alpha)$ shifts signal energy towards directions that carry smaller singular values. This raises $d^{\dagger}$, and consequently, the achievable risk.
This experiment illustrates that ESD, rather than raw spectral decay, is the pivotal measure governing learnability.
@@ 8 Extension to RKHS Regression
This section extends the concepts of Effective Span Dimension (ESD) and span profile, developed in Section˜3, to the setting of RKHS regression. Since our goal is to develop a population level complexity measure, we focus on the simple case where the eigenfunctions of the kernel are fully known and computable, and leave a thorough analysis in future studies.
@@ 8.1 RKHS regression
$y_{i}=f^{*}(x_{i})+\epsilon_{i},\quad\epsilon_{i}\text{ are i.i.d. with }\mathbb{E}[\epsilon_{i}]=0,\ \mathrm{Var}(\epsilon_{i})=\sigma_{0}^{2},\quad i=1,\dots,n,$ (20)
where $x_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mu$, and $f^{*}\in L^{2}(\mathcal{X},\mu)$ is the target function. We use a kernel $\mathbf{k}(\cdot,\cdot)$ with Mercer decomposition $\mathbf{k}(x,x^{\prime})=\sum_{j=1}^{\infty}\lambda_{j}\psi_{j}(x)\psi_{j}(x^{\prime})$, where $\{\psi_{j}\}_{j=1}^{\infty}$ form an orthonormal basis for $L^{2}(\mathcal{X},\mu)$ and $\bm{\lambda}=\{\lambda_{j}\}_{j\geq 1}$ is the sequence of eigenvalues not necessarily sorted. For simplicity, we assume there are no ties among the eigenvalues and let $\pi$ be the permutation that sorts them in descending order, so that $\lambda_{\pi_{1}}>\lambda_{\pi_{2}}>\dots>\lambda_{\pi_{k}}>\ldots$. The coefficients of $f^{*}$ in this basis are $\theta_{j}^{*}=\langle f^{*},\psi_{j}\rangle_{L^{2}(\mu)}$. An estimator $\hat{f}$ has risk $\mathcal{R}(\hat{f};f^{*})=\mathbb{E}\|\hat{f}-f^{*}\|_{L^{2}(\mu)}^{2}$. If we define $\hat{\theta}_{j}=\langle\hat{f},\psi_{j}\rangle_{L^{2}(\mu)}$, we can also write the risk as
@@
Here we kindly remind the reader that in RKHS, we use subscript $i$ for samples and subscript $j$ for eigen-coordinates. The subscript $j$ aligns with the use of notation in the sequence model, where we use indices $j$ to denote coordinates.
@@
Before we introduce the definition of ESD, we demonstrate that the approximation in Equation˜3 can be made exact by increasing the variance to absorb the approximation error.
@@
$=\sum_{k\geq 1}\theta_{k}^{*}\underbrace{\left(\frac{1}{n}\sum_{i=1}^{n}\psi_{k}(x_{i})\psi_{j}(x_{i})\right)}_{:=G_{kj}}\;+\;\underbrace{\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}\,\psi_{j}(x_{i})}_{=\xi_{j}}$
@@
Since $x_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mu$ and $\{\psi_{j}\}$ are orthonormal in $L^{2}(\mathcal{X},\mu)$, we have $\mathbb{E}[G_{kj}]=\mathbf{1}_{\{k=j\}}$. Hence $\mathbb{E}[z_{j}]=\theta_{j}^{*}$.
@@
where we have defined $\Delta_{j}=(G_{jj}-1)\theta_{j}^{*}+\left(\sum_{k\geq 1,k\neq j}G_{kj}\theta_{k}^{*}\right)$. This term does not appear in the sequence model; its randomness arises solely from the random covariates $x_{i}$. As $n\to\infty$, this term vanishes because $G_{jj}=n^{-1}\sum_{i}\psi_{j}\left(x_{i}\right)^{2}\to 1$ and $G_{kj}=n^{-1}\sum_{i=1}^{n}\psi_{j}\left(x_{i}\right)\psi_{j^{\prime}}\left(x_{i}\right)\to 0$ for $k\neq j$. Furthermore, since $\mathbb{E}(\xi_{j}\mid\{x_{i}\}_{i\in[n]})=0$, we have
@@
The presence of $\Delta_{j}$ effectively inflates the variance in $z_{j}$ to $\sigma_{0}^{2}/n+\mathrm{Var}(\Delta_{j})$. One can show that $\mathrm{Var}(\Delta_{j})=\mathrm{Var}(f^{*}(x)\psi_{j}(x))$, which is bounded by $\|f^{*}\|_{\infty}^{2}$. This is how we will control the impact of $\Delta_{j}$ in the following development.
@@ 8.2 ESD for RKHS Regression
We start by analyzing the counterpart of the PC estimator, the Kernel Principal Component Projection Estimator (KPCPE), defined as
$\hat{f}_{k}^{\operatorname{PC}}(x):=\sum_{j:\lambda_{j}\geq\lambda_{\pi_{k}}}z_{j}\psi_{j}(x),$ (22)
@@
The risk $\mathcal{R}_{k}:=\mathbb{E}\|\hat{f}_{k}^{\operatorname{PC}}-f^{*}\|_{L^{2}(\mu)}^{2}$ decomposes into squared bias $B(k)$ and variance $V(k)$. Since $\mathbb{E}[\psi_{j^{\prime}}(x_{i})\psi_{j}(x_{i})]=\mathbf{1}_{\{j=j^{\prime}\}}$, we have $\mathbb{E}[z_{j}]=\theta_{j}^{*}$. Therefore, the bias is due to truncation as
@@
Under Assumption 8.1, $\tau_{j}^{2}\leq\mathbb{E}[f^{*}(x)^{2}\psi_{j}(x)^{2}]\leq\|f^{*}\|_{\infty}^{2}$. Subsequently, the variance is bounded by $V(k)\leq\frac{k}{n}(\sigma_{0}^{2}+\|f^{*}\|_{\infty}^{2})$. This motivates us to define the effective noise variance per component as
@@
We comment that Zhang et al. [50] have established the minimax optimality of the well-tuned PC estimator. Proposition˜8.3 suggests that the risk of the well-tuned PC estimator scales as $d^{\dagger}/n$. Therefore, we essentially express the minimax rate therein using the ESD without reliance on the classical eigen-decay conditions or source conditions.
@@
The kernel $\mathbf{k}$ is said to be $(K,n)$-regular if there are some constants $c_{1}\in(0,1)$ and $C_{1}$ such that $\sum_{i\leq c_{1}K}\lambda_{\pi_{i}}^{-1}\leq C_{1}n$.
@@
$\inf_{\hat{f}}\sup_{f^{*}\in\mathcal{F}_{K,\mathbf{k}}^{(n)}}\mathcal{R}(\hat{f};f^{*})\asymp\frac{\sigma_{0}^{2}K}{n},$ (29)
@@
Combining Theorem˜8.5 and Proposition˜8.3, the optimally tuned KPCPE estimator is minimax rate optimal over $\mathcal{F}_{K,\mathbf{k}}^{(n)}$, with rate $\sigma_{0}^{2}K/n$.
The KPCPE serves as a simple benchmark for spectral methods. This analysis, via the ESD, characterizes the performance of the optimally tuned KPCPE based directly on the properties of the specific signal $f^{*}$ (via $\theta^{*}$) and kernel spectrum $\bm{\lambda}$, without requiring standard assumptions like source conditions or polynomial eigenvalue decay. Therefore, we consider the ESD evaluated at the design-adjusted noise level $\sigma^{2}$ as a key measure of statistical complexity in RKHS regression. In summary, the span profile framework provides a unified perspective on the generalization performance of spectral methods across a variety of models.
@@
Following the framework in Section˜4, we can quantify a class of populations using a quota sequence $\bm{K}=\{K_{n}\}_{n=1}^{\infty}$. For some $n_{0}\in\mathbb{N}_{+}$, define
$\mathcal{F}_{\bm{K},\mathbf{k}}=\{f^{*}\in L^{2}(\mathcal{X},\mu)\cap L^{\infty}(\mathcal{X},\mu):\|f^{*}\|_{\infty}\leq\sigma_{0}C_{0},\quad d^{\dagger}(\bar{\sigma}^{2}/n;f^{*},\mathbf{k})\leq K_{n},\forall n\geq n_{0}\},$ (30)
@@
$\inf_{\hat{f}}\sup_{f^{*}\in\mathcal{F}_{\bm{K},\mathbf{k}}}\mathcal{R}(\hat{f},f^{*}).$ (31)
@@
$\inf_{\hat{f}}\sup_{f^{*}\in\mathcal{F}_{\bm{K},\mathbf{k}}}\mathcal{R}(\hat{f},f^{*})\asymp\bar{\sigma}^{2}\frac{K_{n}}{n}.$
˜8.4 is a mild condition. The following is an example where we use Theorem˜8.6 to recover the minimax convergence rate derived under the classical polynomial eigen-decay conditions and source conditions.
Example 8.7.
Suppose $\mathbf{k}$ admits spectrum such that $\lambda_{\pi_{i}}\asymp i^{-\beta}$ with $\beta>0$. Let $K_{n}=\lfloor n^{\frac{1}{1+s\beta}}\rfloor$ for any $s\geq 1$. It is easy to see that
$\sum_{i\leq c_{1}K_{n}}\lambda_{\pi_{i}}^{-1}\asymp(c_{1}K_{n})^{\beta+1}\asymp n^{\frac{\beta+1}{s\beta+1}}\lesssim n.$
Therefore, the minimax optimal rate for the class $\mathcal{F}_{\bm{K},\mathbf{k}}$ is $\frac{\bar{\sigma}^{2}K_{n}}{n}\asymp\sigma_{0}^{2}n^{-\frac{s\beta}{1+s\beta}}$, which is the same as the optimal rate given by the source condition with smoothness parameter $s$.
Remark 8.8.
In Section˜2, we simplify our discussion by assuming that $\mathbf{k}$ is positive definite. In practice, positive semi-definite (PSD) kernels may also be used. In the case where $\mathbf{k}$ has rank $d<\infty$, spectral algorithms inevitably induce a systematic (squared) bias
$\Delta_{f^{*},\mathbf{k}}=\|f^{*}\|_{L^{2}(\mu)}^{2}-\sum_{j\in[d]}\left(\theta_{\pi_{i}}^{*}\right)^{2},$
regardless the regularization parameter. In that case, we may modify the definition of ESD by adding the systematic bias into the summation to the sum of squared tails.
Specifically, for a PSD kernel $\mathbf{k}$ of rank $d<\infty$, we modify the definition of ESD in Definition˜8.2 as
$\bar{d}^{\dagger}=\bar{d}^{\dagger}(\sigma^{2};f^{*},\mathbf{k})=\min\left\{j\in[d]:\frac{1}{j}\left[\Delta_{f^{*},\mathbf{k}}+\sum_{i=j+1}^{d}\left(\theta_{\pi_{i}}^{*}\right)^{2}\right]\leq\,\sigma^{2}\right\}.$
Again, $\sigma^{2}\bar{d}^{\dagger}$ characterizes the risk of the well-tuned PC estimator as in Proposition˜8.3.